I need help on percent proportions. This problem: 42 is what percent of 48? to find the percantage what do i do ? Help Please !! Thanks Helen

Answer by Here2Help(19)

Hi Helen, I am going to give this a try as this is my first time here.
42/48=x/100
48*x=100*42
48x=4200
48x/48=4200/48
x=87.5%
Hope this is what you needed.